OpenStudy (anonymous):

Find all critical numbers and use the Second Derivative Test to determine all local extrema. y = x^2 - (16)/(x)

OpenStudy (anonymous):

@tcarroll010 can u help?

OpenStudy (anonymous):

are u there?

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

y' = 2x + 16/x^2 , y'' = 2 - 32/x^3

OpenStudy (anonymous):

i'm trying to click on the equation but its not working

OpenStudy (anonymous):

so now i have to set the second derivative = to 0 and solve right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i got x = -1.2599-2.1822 i, x = 2.5198, x = -1.2599+2.1822 i

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what i got is wrong?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so i have correct my derivative right?

OpenStudy (anonymous):

to*

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

idk i dont think i can do this one

OpenStudy (anonymous):

i did this one before it asked the same thing but now the problem was this (x^3 - 3x^2 - 9x) y’ = 3x^2 – 6x – 9 3x^2 – 6x – 9 = 0, x = -1, x = 3 y’’ = 6x – 6 6(-1) – 6 = -12 6(3) – 6 = 12 at -12 it is a maximum and at 12 there is a minimum.

OpenStudy (anonymous):

do i do similar to that or?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yea the numerator is equal to zero

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yea